shift summary for owl shift 06/15/99

Sandrine Laplace (laplace@hep.uchicago.edu)
Tue, 15 Jun 1999 07:22:09 -0500

Shift Summary for owl shift, 06/15/99

On Shift: Sydney, Sandrine

also here: Elizabeth, Rick for one hour.
Arun passed by.

Hours of beam delivered at all intensities: 7
Hours of beam delivered at nominal intensity: 3.5
Hours of E832 data (special runs not included): 7
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:

========================================================================
Runs ended this shift:
13602 - 5.4 million events - BEAM_832

Tapes brought to FCC this shift: 0

========================================================================
Accesses during shift: 0

========================================================================
Overview of shift:

the most quiet shift in the world: Took BEAM_832 data smoothly.
First 3.5 hours - SEM ~1.5E12, BOA ~1.5
Second 3.5 hours - SEM ~4.3E12, BOA ~1.8

========================================================================
Description of Shift:

0:00 Inherit shift with hadrons, run 13602 in progress.
Check crate 16, which was smelling earlier in the evening.
Crate 16 +12V is at 1.8V according to sdaq. +12V is not
used by any module in this crate, but the drop in
voltage is indicative of something wrong here. Should
be investigated more fully on day shift.

0:35 Channel 372 PMT gain drops by 6.5%.

1:10 Elizabeth finds no readout out of MA ADCs.

1:37 Channel 372 gain is now 9% low.

1:43 Channel 372 is back to nominal gain and fine for rest of shift.

2:55 TeV quench at F2, recovery in progress.

3:15 Quench recovered, beam is back at 1.5E12.
Look online histos: PMT gain drift vs DB shift of 1% for MF 2
of CsI. Checked HV, seem OK.

3:22 Arun found that the connector for the cable going from the MA
splitter-shaper to the ADC in crate 16 was misregistered. This
may have been the cause of the earlier problem with this crate.
After spill 187, run 13602, the cable is correctly connected but
the ADC should be swapped on the day shift.

4:15 Intensity up to 4.3E12.

6:45 Beam is shutdown as scheduled.

6:55 Stop run 13602. Start LASER_SCAN 13603.

========================================================================
Follow up Notes on Previous Shift

The Camac crate 16 still smells a little bit, but nothing worse.
Problem may have been due to misregistered cable from the
MA splitter-shape to the MA ADC in this crate.

========================================================================
Items to Follow up on Next Shift

Please watch Camac crate 16 in RR 191 closely. CAMAC crate and/or
power supply may need to be swapped during day shift. MA ADC
definitely should be swapped during day shift.

On the 6/14 evening shift, the L2 dummy busy was set to 50 us
to study DA-related DC noise. This should be set back to
the nominal value.

Regards,

Sandrine and Sydney.